OpenStudy (anonymous):

the polynomial y=(a-2)x^2+(1-3b)x+(5-2c) has at least 3 roots. Find a, b and c.

OpenStudy (anonymous):

Its a quadratic equation and quadratic equation can have maximum 2 roots

OpenStudy (raden):

yea, so that this polynomial has at least 3 roots must be a cubic aquation not quadratic equation

OpenStudy (raden):

general form cubic equation is f(x)=ax^3+bx^2+cx+d

OpenStudy (anonymous):

I'm sure this is a quadratic equation - a zero polynomial I think

OpenStudy (raden):

a zero polynomial to get roots, but ur question is find a,b, and c ? somethink wrong i think, recheck your problem

OpenStudy (anonymous):

Sorry, this is what the question said. Would a zero polynomial have all the coefficients equal to zero? That's what I wrote down but I don't understand it.

OpenStudy (raden):

if that is a quadratic equation, still incorect question :) (becasue quadratic equation for coefficien x^2 cant be equal zero) if no, take : a-2 = 0 a = 2 next, 1-3b = 0 1=3b b=1/3 the last one 5-2c=0 5=2c c=5/2

OpenStudy (anonymous):

hmm ok then, it must be an incorrect question. Thanks anyways! =D

OpenStudy (raden):

welcome :D